If our daughter is 5, and I am 30. How old will we be when our daughter is 3 times my age? I know that I was 25 when she was born. I thought I had it for a second, with her being 15 and me being 40, but then I realized that that was 5 short of being 3x's my age. I am so confused, and my instructor is a no show..Please help me figure out how to do this: )
Answer by jrfrunner(365)

you must mean, how old will you be when you are 3 times your daughters age, not the other way around
--
your age=30, your daughter =5
--
let x=additional years
--
30+x=3*(5+x)
30+x=15+3x
30-15=3x-x
15=2x
7.5=x
---
in 7.5 years you will be 3 times your daughter's age
validate
--
30+x=3*(5+x)
30+7.5=3*(5+7.5)
37.5=3*12.5=37.5